### v1

# Label-Guided In-Context Learning for Named Entity Recognition

Abstract
In-context learning (ICL) enables large language models (LLMs) to perform new tasks using only a few demonstrations. In Named Entity Recognition (NER), demonstrations are typically selected based on semantic similarity to the test instance, ignoring training labels and resulting in suboptimal performance. We introduce Deer, a new method that leverages training labels through token-level statistics to improve ICL performance. Deer first enhances example selection with a label-guided, token-based retriever that prioritizes tokens most informative for entity recognition. It then prompts the LLM to revisit error-prone tokens, which are also identified using label statistics, and make targeted corrections. Evaluated on five NER datasets using four different LLMs, Deer consistently outperforms existing ICL methods and approaches the performance of supervised fine-tuning. Further analysis shows its effectiveness on both seen and unseen entities and its robustness in low-resource settings. [^1] Code and data are available at https://github.com/bflashcp3f/deer.
Label-Guided In-Context Learning for Named Entity Recognition
[…]
1 Introduction
Identifying and categorizing named entities in text (Named Entity Recognition, NER) is a core information extraction (IE) task with applications in diverse domains including science (Lin et al., 2004; Leser and Hakenberg, 2005; Luan et al., 2018; Bai et al., 2022), news (Sang and De Meulder, 2003; Whitelaw et al., 2008), and social media (Finin et al., 2010; Ritter et al., 2011; Li et al., 2012). Supervised learning works well for NER using models with tailored architectures, fine-tuned on large labeled datasets (Nadeau and Sekine, 2007; Ma and Hovy, 2016). However, these approaches do not generalize well across domains or to new entity types without re-training on annotated data, thereby constraining their broader applicability (Lin and Lu, 2018; Bai et al., 2021).
Figure 1: Unlike existing in-context learning methods for NER that select demonstrations solely based on input embedding similarity, our proposed method, Deer, leverages token-level, label-related statistics to enhance demonstration retrieval in a training-free manner. It further uses this information to reflect on initial predictions and make refinements.
Figure 2: Overview of Deer. In the preparation stage (Step 0), the method compiles training input and labels to compute token frequencies and probabilities in three scenarios: 1) entity token, 2) context token, and 3) other token, along with their associated spans. In the inference stage (Step 1 and 2), Step 1 retrieves sentence-level demonstrations by emphasizing potential entity- and context-related tokens based on probabilities from Step 0. Step 2 refines predictions from Step 1 by addressing error-prone tokens based on label statistics, focusing on three token types: unseen tokens, "false negative" tokens, and boundary tokens. For each token type, the refinement process retrieves span-level demonstrations and prompts LLMs to adjust predictions. See §2 for further details.
Recent advances in large language models (LLMs; Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; OpenAI et al., 2024) suggest an alternative: in-context learning (ICL; Dong et al., 2024), where models learn new tasks directly from examples provided as input, without parameter updates. As shown in Figure 1, a common ICL strategy for NER (Wang et al., 2023a; Lu et al., 2024) involves retrieving a fixed number of demonstrations using embeddings of input data, and then prompting the model to extract entities. While such methods show particular promise in low-resource scenarios (Lee et al., 2022), it does not scale as efficiently as supervised methods (Sainz et al., 2024). For example, Monajatipoor et al. (2024) report a 79.3 F1 score on the NCBI dataset using GPT-4 with 16 examples per query. In contrast, a fine-tuned BERT-base model achieves an 84.0 F1 score when given access to the same training data. This discrepancy suggests that the current ICL approach makes less effective use of the training data. For demonstration retrieval, similarity between the test example and training examples depends on sentence-level semantic similarity and does not use the label on the training instances. Task-agnostic sentence embeddings fail to capture token-level label details (e.g., entities), which are critical for NER. Similarly, the selected examples included in the prompt are the totality of label information provided to the model; the vast majority of annotations are ignored. As a result, the model’s predictions may miss known entities, invent spurious ones, or have inaccurate boundaries.
Motivated by these findings, we propose Deer: Data statistics-grounded namEd Entity Recognition, an ICL method that leverages label statistics to support collective learning from the entire training set. Deer operates in two steps. First, it improves demonstration retrieval by aligning it with the token-level nature of NER. Rather than relying on task-agnostic sentence embeddings, it uses token-based retrieval guided by label statistics to prioritize tokens likely to belong to entities or their context, resulting in more relevant examples. Second, it introduces an error reflection step that identifies potentially misclassified tokens using the same token-level label statistics. These tokens are revisited with targeted span-level demonstrations, allowing the model to refine its predictions and improve overall performance.
We evaluate Deer on five NER datasets spanning four domains: news, biomedicine, social media, and mixed web text, using four recent LLMs: Qwen2.5-7B, Llama3.3-70B, GPT-4o-mini, and GPT-4o. The results show that Deer consistently outperforms ICL baselines across datasets and models. With GPT-4o, Deer achieves performance competitive with supervised models on four out of five datasets, while the open-source Llama3.3-70B also demonstrates strong results. Further analysis confirms that Deer enhances example retrieval, improves recognition of both seen and unseen entities, and demonstrates robustness in low-resource settings.
2 Deer [Uncaptioned image]
We introduce Deer, an ICL approach designed to collectively leverage the full training labels while focusing on token-level details for NER. Deer comprises two steps, as illustrated in Figure 2: 1) label-guided retrieval, and 2) error reflection. The following sections detail the problem setup and our training data preparation process for Deer, followed by a description of each of its two steps.
2.1 Problem Setup and Training Data Preparation
Named entity recognition aims to identify all text spans corresponding to pre-defined entity types in a sentence. Following prior work on LLM-based entity recognition (Dunn et al., 2022; Lu et al., 2022), we formulate the task as a span identification problem and format the output as a structured JSON object (Figure 1), using the key “named_entities” for a list of JSON objects, each containing the name and type of an entity (e.g., {“name”: “Barack Obama”, “type”: “PERSON”}). In Appendix B, we compare it with two other prompting formats from recent work and find that our chosen format yields the best performance.
We support Deer through a pre-processing step: we categorize training input tokens into three distinct classes based on entity labels and compute the frequency and probability of each token within these classes (Step 0 in Figure 2). Entity tokens are those that belong to an identified entity. Context tokens are the tokens surrounding an entity, typically two tokens on each side, which provide identification cues and help define entity boundaries (Huang et al., 2015). Other tokens include all remaining tokens in the sentence. For example, consider the sentence in Figure 2, ‘‘The risk of cancer, ...’’, where “cancer” is an entity token, “risk” and “of” are context tokens, and “The” is an other token, as it falls out of the context window. Context tokens like “risk of” provide contextual clues for identifying Disease entities, even in cases where entities are unseen in the training data.
Moreover, we define token-focused spans for these categories: entity spans include the entity along with its surrounding context, context spans are similar but correspond to context tokens, and other spans consist of two adjacent tokens on each side of an other token. For instance, the span ‘‘risk of cancer, especially’’ serves as an entity span for “cancer” but as a context span for “risk.” [^2] In our implementation, entity spans exclude neighboring entities within the context window to encourage greater attention on context tokens for more generalizable entity recognition. For example, given the sentence “Lionel Messi and Cristiano Ronaldo are exceptional football players” and a context window of two, the entity spans for the two PERSON entities are ‘‘Lionel Messi and’’ and ‘‘and Cristiano Ronaldo are exceptional’’. These label-dependent spans offer a more precise and structured approach to handling tokens in different scenarios. In §4.1, we analyze the impact of incorporating these three token types and demonstrate the importance of context tokens, particularly for improving performance on unseen entities.
2.2 Label-guided Retrieval
Deer begins with label-guided, token-based retrieval for ICL (Step 1 in Figure 2). The ICL prompt consists of the task description, specifying the targeted entity types and output format, followed by $N$ examples selected from the training set and the query sentence. Unlike previous work that uses task-agnostic sentence embeddings, we aim to address the token-centric nature of NER. Specifically, to compute the similarity between a training sentence $s_{i}$ and a test sentence $s_{t}$, we define a similarity score $\mathbf{S}_{s_{i},s_{t}}$ that combines exact token matching and embedding-based semantic similarity:
$\mathbf{S}_{s_{i},s_{t}}=\lambda_{1}\cdot S_{\text{token}}(s_{i},s_{t})+\lambda_{2}\cdot S_{\text{embed}}(s_{i},s_{t}),$ (1)
where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters controlling the relative contribution of two components. The token match score $S_{\text{token}}$ is computed as:
$S_{\text{token}}(s_{i},s_{t})=\sum_{t\in s_{t}}\mathbb{I}(t\in s_{i})\cdot W(t),$ (2)
where $\mathbb{I}(t\in s_{i})$ is an indicator function that equals 1 if test token $t$ appears in the training sentence $s_{i}$, and 0 otherwise. The function $W(t)$ assigns importance to each token:
$W(t)=\begin{cases}w_{e}P(t_{e})+w_{c}P(t_{c})+w_{o}P(t_{o}),&\text{if }t\text{is seen},\\ 1,&\text{otherwise }\end{cases}$ (3)
where $P(t_{e})$, $P(t_{c})$, and $P(t_{o})$ are the probabilities of token $t$ being an entity, context, or other token, respectively, in the training corpus. The hyperparameters $w_{e}$, $w_{c}$, and $w_{o}$ control the relative importance of these token types, allowing the model to prioritize entity and context tokens, which are more informative for NER. [^3] For all hyperparameters, we perform a grid search over three possible values: $[1.0,0.5,0.01]$, and find that setting $\lambda_{1}=1.0$, $\lambda_{2}=1.0$, $w_{e}=1.0$, $w_{c}=1.0$, and $w_{o}=0.01$ performs well in most cases. See ablations in Appendix B. For semantic similarity, we compute a sentence embedding by aggregating token embeddings:
$\mathbf{v}_{s}=\sum_{t\in s}W(t)\cdot\mathbf{v}_{t},$ (4)
where $\mathbf{v}_{t}$ represents the embedding of token $t$. The final score is computed via cosine similarity:
$S_{\text{embed}}(s_{i},s_{t})=\cos(\mathbf{v}_{s_{i}},\mathbf{v}_{s_{t}}).$ (5)
By integrating both token matching and embedding-based similarity, our method ensures precise retrieval of seen tokens while also accounting for unseen tokens, leading to optimized performance across different dataset characteristics.
2.3 Error Reflection (ER)
Even with improved retrieval, ICL can only incorporate a small subset of training examples, leaving blind spots that limit the model’s ability to make well-informed predictions. For example, the model may miss entities present in the training set but absent from retrieved examples or make boundary errors due to insufficient guidance. To address this, we introduce an error reflection step that uses training statistics to identify error-prone tokens and provides targeted demonstrations to help the model revise its predictions. We focus on single-entity, boundary-related errors, which have been identified as a major source of NER mistakes (Ding et al., 2024; Lu et al., 2024). A hierarchical breakdown of NER errors is shown in Figure 4 in the appendix.
As shown in Step 2 of Figure 2, our reflection step targets three token categories associated with potential errors: unseen tokens, “false negative” tokens, and boundary tokens of predicted entities. The first two categories help recover missed entities, while the third addresses both false positives and boundary-related errors. The reflection process proceeds sequentially. We first examine unseen and false negative tokens in the query sentence to identify additional candidate entities. Then, we refine or discard predictions by inspecting boundary tokens. To enable this process, we use a chain-of-thought reasoning approach (Wei et al., 2023), prompting the model to reason about the relationship between the query and retrieved demonstrations before making final predictions. Details on how we identify these tokens and construct reflection prompts for each category are provided below. Exact prompts are listed in Appendix A.1.
Unseen Tokens
Unseen tokens, defined as tokens not present in the training data, often lead to false negatives since they lack direct matches for retrieval. To address this, we leverage their surrounding context. Specifically, we select unseen tokens that were not predicted as part of an entity but are surrounded by tokens frequently labeled as entities or context tokens in the training set. This targeted strategy avoids reflecting on every unseen token, reducing inference cost while improving prediction accuracy (see Table 11 in the appendix for reflection frequency statistics). For each selected unseen token, we retrieve $M$ span-level demonstrations from the training set by matching its context tokens. A chain-of-thought prompt then guides the model to reason through these demonstrations and determine whether a new entity should be extracted. For sentences with multiple unseen tokens, this process is applied in the order the tokens appear.
“False Negative" Tokens
These are tokens commonly annotated as entities in the training set but missed during initial predictions. A token is flagged as a “false negative" if its entity likelihood in the training set exceeds a predefined threshold. For each such token, we retrieve $M$ span-level examples containing it and prompt the model to predict whether an entity should be extracted.
Boundary Tokens
These tokens lie at the edges of predicted entities and are critical for refining entity spans or eliminating false positives. For each prediction, we examine $2K$ tokens around the boundary: $K$ edge tokens of the entity and $K$ from the surrounding context. We focus on tokens likely to be misclassified based on training statistics, for example, tokens predicted as part of an entity but appeared more as context tokens in the training set. For each reflected token, we retrieve span-level examples from all three categories: entity, context, and other. This variety allows the model to reason through different scenarios and decide whether to include each token, ultimately refining the entity’s boundaries.
3 Experimental Setup
3.1 Dataset and Evaluation Metric
To demonstrate the effectiveness of our method, we evaluate it on five NER datasets spanning four domains, which have been used in recent ICL-based NER studies (Wang et al., 2023a; Monajatipoor et al., 2024). These datasets include CoNLL03 (Sang and De Meulder, 2003) for news, NCBI-disease (Doğan et al., 2014) and bc2gm (Smith et al., 2008) for biomedicine, TweetNER7 (Ushio et al., 2022) for social media, and OntoNotes (Weischedel et al., 2013) for web text from various sources. Dataset statistics are provided in Table 1. We refer readers to the corresponding papers for details on each dataset. Following prior work (Berger et al., 2024; Bai et al., 2024), we control computational costs by limiting the test set size to 1,000 examples through random sampling of their original test sets. To ensure the sampled subset provides a reliable estimate of model performance, we calculate 95% confidence intervals using 1,000 bootstrap resamples, and the margins of error fall within a 0.01-0.03 range, verifying the validity of our performance estimates. Further details on this process can be found in Appendix A.2.
All experiments are evaluated using the standard mention-level matching metric for NER (Sang and De Meulder, 2003), where a predicted entity is considered correct only if it matches both the span boundary and the entity type.
# Sentences
Dataset Domain # Ent. (Tr / Dv / Ts)
CoNLL03 (2003) News 4 14.0k / 3.3k / 3.5k
OntoNotes (2013) Web 18 60.0k / 8.5k / 8.3k
NCBI (2014) Biomed 1 5.4k / 0.9k / 0.9k
BC2GM (2008) Biomed 1 12.5k / 2.5k / 5.0k
TweetNER7 (2022) Social 7 7.1k / 0.8k / 0.6k
Table 1: Statistics of the five experimented datasets.
LLM Method NCBI bc2gm CoNLL03 OntoNotes TwNER7 average
Qwen2.5-7B BM25 66.8 57.0 83.0 62.6 55.0 64.9
BSR (2023) 68.0 59.9 86.0 66.8 56.1 67.4
Kate (2022) 66.9 59.3 83.7 63.8 57.1 66.2
DEERw/o ER 71.7 62.5 87.2 68.4 57.3 69.4
DEER 73.2 63.6 87.8 71.8 57.5 70.8
Llama3.3-70B BM25 79.7 65.5 88.9 76.0 61.7 74.4
BSR (2023) 79.3 67.7 90.2 78.9 62.8 75.8
Kate (2022) 79.4 66.5 88.3 76.1 62.0 74.5
DEERw/o ER 81.9 68.6 89.6 78.9 63.3 76.5
DEER 83.9 69.6 90.2 80.2 63.5 77.5
GPT-4o-mini BM25 74.1 62.9 87.8 74.4 59.3 71.7
BSR (2023) 74.9 62.9 89.6 77.4 61.1 73.2
Kate (2022) 76.0 64.1 89.2 74.9 59.6 72.8
DEERw/o ER 77.6 65.4 89.9 77.0 60.4 74.1
DEER 80.3 67.1 90.9 79.2 61.6 75.8
GPT-4o BM25 81.0 71.9 92.0 81.2 62.8 77.8
BSR (2023) 81.2 73.5 93.3 82.9 63.9 78.9
Kate (2022) 80.0 73.6 92.9 80.7 62.2 77.9
DEERw/o ER 82.7 74.3 93.5 83.5 63.6 79.5
DEER 84.8 75.4 93.6 85.2 64.3 80.7
Table 2: Comparison of Deer with three ICL baselines across five datasets using four recent LLMs and 8 demonstrations per query. Deerw/o ER excludes the error reflection step and already outperforms all baselines, validating the effectiveness of our label-guided retrieval. Adding error reflection yields further improvements. Among the LLMs, GPT-4o performs best overall, while Llama3.3-70B surpasses GPT-4o-mini and approaches GPT-4o, demonstrating the promise of open-source models.
Unseen Entity
Seen Entity Seen Token Unseen Token
Kate Deer Kate Deer Kate Deer
NCBI 88.7 91.4 61.1 66.5 70.3 81.0
bc2gm 85.8 88.4 60.5 60.1 70.8 75.0
CoNLL03 95.5 95.7 84.9 87.7 91.2 92.1
OntoNotes 86.0 89.6 65.5 72.4 76.1 79.7
TweetNER7 71.1 74.9 45.9 46.9 58.3 58.3
Table 3: Breakdown of test set performance for Kate and Deer (with GPT-4o) based on entity presence in the training data. Unseen entities are further categorized by whether they contain novel tokens. Deer outperforms Kate on both seen and unseen entities, highlighting its effectiveness to novel entities.
8 demo. 32 demo. Fine-tuned
Dataset Deerw/o ER Deer Deerw/o ER Deer BERT UniNER
NCBI 82.7 (\$2.1) 84.8 (\$3.0) 84.1 (\$6.5) 85.7 (\$7.3) 84.0 87.0
bc2gm 74.3 (\$2.9) 75.4 (\$3.7) 75.0 (\$9.4) 75.6 (\$9.9) 79.0 82.4
CoNLL03 93.5 (\$1.9) 93.6 (\$2.2) 93.9 (\$5.8) 94.0 (\$6.0) 90.4 93.3
OntoNotes 83.5 (\$2.5) 85.2 (\$3.3) 85.8 (\$7.5) 86.6 (\$8.2) 87.2 89.9
TweetNER7 63.6 (\$1.9) 64.3 (\$3.4) 64.9 (\$6.1) 65.3 (\$7.2) 60.9 65.7
Table 4: ICL results of Deer (using GPT-4o) with 8 and 32 demonstrations, compared to two fine-tuning baselines. Values in parentheses indicate associated API inference costs on the test set. Increasing the number of demonstrations improves performance, and error reflection continues to provide additional gains at larger scales. Notably, Deer with 8-shot error reflection performs competitively with the 32-shot setting without reflection, but at a significantly lower cost. Deer also matches or exceeds BERT-base on most datasets and narrows the gap with UniNER.
3.2 Baselines
We compare Deer against both ICL and fine-tuning baselines to demonstrate its improvements over standard ICL methods and its ability to close the performance gap with fine-tuning. All methods use the full training set for either fine-tuning or in-context demonstration selection.
ICL Baselines
We include three ICL baselines: KATE (Liu et al., 2022), BM25 (Robertson et al., 2009), and BSR (Gupta et al., 2023). Kate serves as our primary ICL baseline, as it remains the dominant framework for ICL-based NER (Wang et al., 2023a; Berger et al., 2024; Monajatipoor et al., 2024). This method selects in-context examples through nearest neighbor search using sentence-level embeddings. BM25 treats both the test input and demonstrations as bags of words and measures their relevance via a TF-IDF–weighted recall of overlapping words. Recent IE work (Sun et al., 2024) has shown its effectiveness for demonstration retrieval. BSR employs BERTScore-Recall (Zhang et al., 2020) as a similarity metric, which measures token-level recall using token embeddings. While mainly used in semantic parsing, we adapt it here for NER. To ensure the competitiveness of the baselines, we evaluate seven embedding models commonly used in recent studies (see Appendix B). OpenAI’s text-embedding-3-small consistently outperforms open-source alternatives such as all-mpnet-base-v2, and is therefore used in all our experiments.
Fine-tuning Baselines
We include two supervised baselines: BERT fine-tuning and UniversalNER-7B (Zhou et al., 2024). We fine-tune a BERT-base model (Devlin et al., 2019) for token-level classification with three random seeds and report the average performance. UniversalNER explores targeted distillation for NER, leveraging gpt-3.5-turbo to generate entity annotations for unlabeled text. The refined dataset is then used to train a student model, which is further fine-tuned on supervised NER datasets. This method achieves competitive results with state-of-the-art NER-specific LLMs, including InstructUIE (Wang et al., 2023b) and GoLLIE (Sainz et al., 2024). [^4] As noted in its GitHub repository, the authors have not released the code and prompts required to reproduce most of the supervised dataset results due to licensing restrictions. Therefore, we rely on the reported results from the original paper for comparison.
3.3 Implementation Details
We conduct experiments with four recent LLMs, covering both open-source and proprietary models: Qwen2.5-7B (Qwen et al., 2025), Llama3.3-70B (Dubey et al., 2024), GPT-4o-mini, and GPT-4o (OpenAI et al., 2024). [^5] In our preliminary experiments, we also evaluated Llama3.1-8B. However, it failed to generate outputs in the required format during ICL (see Appendix B.6 for more details). Thus, we did not pursue it further. We use the same prompt across four LLMs and set the decoding temperature to 0 to facilitate result reproducibility. For main ICL experiments, we retrieve eight demonstrations per query, sorted in ascending order of similarity to the test example. Appendix A contains detailed descriptions of the prompts and all hyperparameters.
4 Results
Comparison with ICL Baselines
Table 2 presents the results of three ICL baselines and two variants of Deer, including the one without error reflection (Deerw/o ER), evaluated across five datasets and four LLMs, using eight retrieved demonstrations per query. Deerw/o ER consistently outperforms Kate in all settings, demonstrating the effectiveness of our label-guided, token-level demonstration retrieval. It also performs better than BM25 and BSR. Interestingly, BSR consistently outperforms Kate, suggesting that BERTScore-Recall, which emphasizes token-level coverage, may be more suitable than cosine similarity on sentence representations for token-centric tasks like NER. A promising future direction would be to integrate our label-guided retrieval with BERTScore-Recall. Adding error reflection leads to further gains across all settings, confirming its effectiveness. A breakdown of performance across the three reflection steps is provided in Table 7 in the appendix. Appendix C (Table 18) presents a detailed error analysis, highlighting Deer’s advantages in reducing under-predicted entities and boundary errors.
Among the LLMs, GPT-4o achieves the best overall results while our method delivers larger relative improvements on smaller models, such as Qwen2.5-7B. For example, on OntoNotes, Deer achieves an 8.0 absolute F1 gain over Kate. Llama3.3-70B shows strong performance, outperforming GPT-4o-mini and approaching GPT-4o, highlighting the potential of open-source LLMs.
Table 3 further analyzes performance based on entity presence in the training data. The test set entities are categorized as either seen or unseen during training. Unseen entities are further divided into those containing novel tokens and those formed by permutations of seen tokens. The results show that Deer consistently outperforms Kate on both seen and unseen entities, demonstrating its effectiveness in generalizing to novel entities.
Scaling Demonstrations and Comparison with Fine-tuning Baselines
To evaluate the scalability of Deer and its performance relative to fine-tuning, Table 4 reports results using GPT-4o with 8 and 32 demonstrations per query, along with two fine-tuning baselines. We include both performance and API inference costs, with and without error reflection. [^6] Inference costs are calculated using OpenAI’s official pricing: \$2.50 per million input tokens and \$10.00 per million output tokens. Increasing the number of demonstrations to 32 consistently improves performance across all datasets. Error reflection continues to provide added benefit even at this larger scale. Notably, Deer with error reflection and 8 demonstrations performs competitively with the 32-demo version without reflection, but at significantly lower cost. This highlights a potential cost-performance trade-off between scaling demonstrations and applying error reflection, which we leave for future work.
Compared to fine-tuning baselines, Deer matches or exceeds BERT-base on four of five datasets and narrows the performance gap with UniNER. These results suggest that reliance on task-specific NER models may decline over time, especially as general-purpose LLMs become more cost-efficient and can be reused across multiple tasks, while specialized models incur additional maintenance and deployment costs.
4.1 Ablation Studies & Analyses
Label-guided Retriever
To better understand the effectiveness of our retriever, we compare three retrieval methods across all five datasets using GPT-4o: sentence embeddings from Kate, our proposed token-based retriever, and an unweighted variant that excludes label-guided token statistics. All embeddings are generated using OpenAI’s text-embedding-3-small model. As shown in Table 5, our label-guided method consistently outperforms the other two. The unweighted variant performs similarly to or worse than sentence embeddings, underscoring the importance of incorporating label-aware, token-level weighting in retrieval. Further analysis of the two components in our retriever is provided in Appendix B.
Retriever Sentence Token (uwg.) Token (wg, ours)
NCBI 80.0 80.2 82.7
bc2gm 73.6 71.6 74.3
CoNLL03 93.1 92.6 93.5
OntoNotes 80.7 81.8 83.5
TweetNER7 62.2 62.7 63.6
Table 5: Comparison of three demonstration retrievers: sentence embeddings (Kate), our label-guided token-based retriever, and an unweighted variant without label statistics. Embeddings are generated using OpenAI’s text-embedding-3-small. Our method consistently outperforms the others, highlighting the value of label-aware, token-level weighting.
Token Types & Context Length
A key contribution of Deer is the introduction of three token types, particularly context tokens, defined as tokens surrounding annotated entities within a specified context window. We conduct an ablation study on different context lengths, ranging from 0 to 3, where 0 indicates the removal of context tokens, leaving tokens classified only as entity or other tokens. The ICL results of Deer on three datasets, presented in Figure 3, show that incorporating context tokens significantly boosts performance, although the improvement plateaus once the context length reaches 2. Moreover, Table 8 in the appendix provides a breakdown of performance on seen and unseen entities as in Table 3, confirming that context tokens are effective in improving performance on unseen entities, supporting our hypothesis.
Figure 3: Comparison of different context lengths in Deer across three datasets, where 0 indicates that tokens are classified as either entity or other tokens. The substantial performance gain from context length 1 over 0 highlights the effectiveness of incorporating context tokens in our approach.
Low-Resource Results
Thus far, our evaluation has focused on high-resource settings, where training sets contain at least 5,000 examples. To assess Deer’s effectiveness in low-resource scenarios, we downsample NCBI to three smaller scales: 16, 64, and 256 examples, and compare its performance against Kate. The results in Table 6 show that Deer consistently outperforms Kate, demonstrating its robustness and effectiveness in low-resource conditions.
Method 16-shot 64-shot 256-shot
Kate 70.3 73.2 73.6
Deer 75.7 76.0 78.1
Table 6: Low-resource experiments on NCBI. We compare Deer with Kate (using GPT-4o) across three scales: 16, 64, and 256 examples. Deer surpasses Kate in all settings, highlighting its effectiveness in low-resource settings.
5 Related Work
LLM-Based Named Entity Recognition
The rise of LLMs has opened new avenues for NER by leveraging their capabilities in instruction-following (Ouyang et al., 2022) and in-context learning (Brown et al., 2020). Recent work on LLM-based NER follows two main directions. One focuses on training NER-specialized models by fine-tuning open-source LLMs (Wang et al., 2023b; Zhou et al., 2024; Sainz et al., 2024). The other, which we follow, explores adapting general-purpose LLMs for NER via prompting (Xie et al., 2023; Ashok and Lipton, 2023; Han et al., 2024). In this direction, the dominant framework, Kate (Liu et al., 2022), selects demonstrations using sentence embeddings. Prior research mainly varies prompting templates (Li et al., 2023) or embedding models for retrieval (Monajatipoor et al., 2024), without addressing the core limitations of Kate in leveraging labeled data for token-level tasks. We address this gap by introducing a label-guided retriever and an error reflection strategy. The most relevant prior work is Wang et al. (2023a), which proposes an entity-based retriever that requires training a BERT-based NER model to guide ICL retrieval. However, the resulting ICL model underperforms the BERT-based retriever itself, limiting its practical value. In contrast, our method shows that general-purpose LLMs can achieve competitive performance with supervised models in a training-free manner.
Reflection Using LLMs
An emerging capability of LLMs is their ability to reason (Wei et al., 2023) and self-reflect (Asai et al., 2023) on their predictions, enabling them to correct errors. Recent research has explored this capability across a wide range of tasks, like code generation (Madaan et al., 2023) and planning (Shinn et al., 2023). Similar approaches have also been applied to NER (Wang et al., 2023a; Chen et al., 2024). These works primarily rely on LLMs’ internal knowledge for decision-making. In contrast, our approach grounds model predictions to the label-based training statistics. This grounding enables the model to capture fine-grained dataset-specific artifacts, akin to supervised methods, and helps close the performance gap between prompting-based and supervised approaches.
6 Conclusion
We present Deer, an ICL-based NER method that enhances demonstration selection through token-level, label-guided retrieval and improves prediction through targeted error reflection. Evaluated on five NER datasets with four LLMs, Deer consistently outperforms existing ICL methods and achieves performance comparable to supervised fine-tuning. Further analyses demonstrate its effectiveness on both seen and unseen entities and its robustness in low-resource settings.
Limitations
This paper focuses on flat NER while acknowledging that other NER challenges, such as discontinuous NER (Dai et al., 2020) and nested NER (Finkel and Manning, 2009), pose unique difficulties, particularly for autoregressive LLMs. Exploring these more complex NER tasks is left for future work. Our method primarily targets demonstration selection within the in-context learning framework. Future research could extend this by jointly optimizing both demonstration selection and task instructions (Pang et al., 2023), as incorporating richer task-specific prompts may improve generalization in more difficult scenarios. Finally, the error reflection mechanism used in this work is guided by domain knowledge and limited to three predefined error types. Recent studies have shown that LLMs are capable of detecting their own errors (Wang et al., 2024; Kamoi et al., 2024), suggesting the potential for automating error reflection using more fine-grained or adaptive error categories.
Appendix A Implemetation Details
A.1 Deer Prompts
We use the CoNLL03 dataset (Sang and De Meulder, 2003), which targets the identification of "Person," "Organization," "Location," and "Miscellaneous" entities, as an example to illustrate the prompts used in our experiments.
In-Context Learning - Input
Here is the JSON template for named entity recognition:
{"named entities": [{"name": "ent_name_1", "type": "ent_type_1"}, ..., {"name": "ent_name_n", "type": "ent_type_n"}]}
Please identify the four types of named entities: "PER", "LOC", "ORG", and "MISC" ("Miscellaneous"), following the JSON template listed above, and output the JSON object. If no named entities identified, output {"named entities": []}.
Input: The girl , who was accompanied to Philadelphia by her parents , will need more surgery later to correct the condition on her chest , back and legs , the hospital said .
Output: {"named entities": [{"name": "Philadelphia", "type": "LOC"}]}
Input: " I know what I m here for , " said Medvedev , who lost in the second round of the Open the last two years after reaching the quarters in 1993 , the same year he tried his hand as a restaurant critic .
Output: {"named entities": [{"name": "Medvedev", "type": "PER"}, {"name": "Open", "type": "MISC"}]}
Input: The church in Australia said on Monday Lynch , Batchelor , Barton and Riel were held in a prison until the weekend , when they were moved to join the other captives at the compound .
Output: {"named entities": [{"name": "Australia", "type": "LOC"}, {"name": "Lynch", "type": "PER"}, {"name": "Batchelor", "type": "PER"}, {"name": "Barton", "type": "PER"}, {"name": "Riel", "type": "LOC"}]}
Input: In a telephone call to a local newspaper from his holiday home in Spain , Dalglish said : " We came to the same opinion , albeit the club came to it a little bit earlier than me . "
Output: {"named entities": [{"name": "Spain", "type": "LOC"}, {"name": "Dalglish", "type": "PER"}]}
Input: Bosnian refugees in Hungary , the first to vote last weekend in their country s first post-war election , found the rules confusing and some had no idea who they voted for , refugees and officials said on Wednesday .
Output: {"named entities": [{"name": "Bosnian", "type": "MISC"}, {"name": "Hungary", "type": "LOC"}]}
Input: Glasgow Rangers striker Ally McCoist , another man in form after two hat-tricks in four days , was also named for the August 31 World Cup qualifier against Austria in Vienna .
Output: {"named entities": [{"name": "Glasgow Rangers", "type": "ORG"}, {"name": "Ally McCoist", "type": "PER"}, {"name": "World Cup", "type": "MISC"}, {"name": "Austria", "type": "LOC"}, {"name": "Vienna", "type": "LOC"}]}
Input: Austrian television said the coach , which was carrying 45 , was en route from the Czech Republic to Italy when the accident occurred near Steinberg , 200 km southwest of Vienna .
Output: {"named entities": [{"name": "Austrian", "type": "MISC"}, {"name": "Czech Republic", "type": "LOC"}, {"name": "Italy", "type": "LOC"}, {"name": "Steinberg", "type": "LOC"}, {"name": "Vienna", "type": "LOC"}]}
Input: Austrian television reported earlier that more than 20 had been hurt in the accident at the station in Linz , 300 km ( 180 miles ) west of Vienna .
Output: {"named entities": [{"name": "Austrian", "type": "MISC"}, {"name": "Linz", "type": "LOC"}, {"name": "Vienna", "type": "LOC"}]}
Input: The fans , in Austria to watch their team play Rapid Vienna last Wednesday , may have been involved in a pub brawl earlier , the spokeswoman said .
Output:
In-Context Learning - Output
{"named entities": [{"name": "Austria", "type": "LOC"}, {"name": "Rapid Vienna", "type": "ORG"}]}
Reflect Unseen - Input
<input_text>
Bitar pulled off fine saves whenever they did .
</input_text>
<candidate_tokens>
[’Bitar’]
</candidate_tokens>
<candidate_token>
Bitar
</candidate_token>
<potential_context_tokens_around>
[’pulled’]
</potential_context_tokens_around>
<context_token>
pulled
</context_token>
<examples>
Input: 10th-ranked American Chanda Rubin has pulled
Output: {"name": "Chanda Rubin", "type": "PER"}
</examples>
Given the input text from a Reuters news article, evaluate each candidate token along with the surrounding tokens to determine if it should be categorized as (part of) one of the four named entity types: "PER" (Person), "LOC" (Location), "ORG" (Organization), or "MISC" (Miscellaneous). **Use provided examples, if available, for reference.** If it should be a new entity, extract the exact text span in the sentence, including any spaces, in JSON format. Ensure tokens are not split (e.g., maintain hyphenated words intact). Note that abbreviations and full names are separate entities. If no changes are made, return {"named entities": []}.
JSON Template:
{"named entities": [{"name": "ent_name_1", "type": "ent_type_1"}, ..., {"name": "ent_name_n", "type": "ent_type_n"}]}
Output Format:
Candidate Token: token_1
Contextual Meaning: ...
Relation to Examples Provided: ...
Rationale: ...
Updates: ... (add a new entity/no change)
Candidate Token: token_n
Contextual Meaning: ...
Relation to Examples Provided: ...
Rationale: ...
Updates: ... (add a new entity/no change)
Final predicted entities for the input text (JSON format):
Reflect Unseen - Output
Candidate Token: Bitar
Contextual Meaning: "Bitar" appears to be a proper noun, likely referring to a person involved in a sporting context, as indicated by the phrase "pulled off fine saves."
Relation to Examples Provided: The example provided features a person (Chanda Rubin) and categorizes her as "PER." This aligns with the context of "Bitar," suggesting it is also a persons name.
Rationale: Given that "Bitar" is a proper noun and is likely a persons name in a sports context, it should be categorized as a person entity.
Updates: Add a new entity.
Final predicted entities for the input text (JSON format):
{"named entities": [{"name": "Bitar", "type": "PER"}]}
Reflect False Negative - Input
<input_text>
Italian Serie A games to be played on Sunday ( league positions in parentheses , all kick- off times GMT ) :
</input_text>
<candidate_tokens>
[’Italian’]
</candidate_tokens>
<candidate_token>
Italian
</candidate_token>
<token_stat>
{"num_occurrences_as_entity": 35, "num_occurrences_as_context_tokens": 0, "num_occurrences_as_other_tokens": 0, "entity_vs_context_count": "35 vs 0", "entity_vs_non_entity_count": "35 vs 0"}
</token_stat>
<examples>
Positive Examples (part of entity):
Input: Middlesbrough s Italian striker Fabrizio
Output: {"name": "Italian", "type": "MISC"}
Hard Negative Examples (context tokens):
Negative Examples (other tokens, not entity nor context):
</examples>
Please follow the instructions below:
1. Evaluate each candidate token listed above to determine if it should be categorized as (part of) one of the four named entity types: "PER" (Person), "LOC" (Location), "ORG" (Organization), or "MISC" (Miscellaneous). Consider both the positive and negative examples provided carefully. Pay particular attention to the overall statistical data on whether tokens are included or excluded from the entity. Hard negative examples highlight tokens that are not part of the entity but are located near it.
2. Review both sets carefully. In many cases, a token may be identified as part of an entity in positive examples but not in negative ones, likely due to inconsistencies in the annotation process. if positive and (hard) negative examples seem similar, **base your decision on statistical data**, such as the frequency of the span being recognized as an entity versus its context, particularly when the data is clear-cut (e.g., one frequency is significantly higher).
3. If the token has not been seen or is rarely seen in the training data, use your best judgment to determine whether it should be considered as part of or the entire name of a **specific** entity.
4. If any modifications are necessary, provide the updated entities by **extracting the exact text span in the sentence**, including any spaces and no outside tokens added, in JSON format. Ensure tokens are not split (e.g., maintain hyphenated words intact). Note that abbreviations and full names are separate entities. If no changes are required, return {"named entities": []}.
JSON Template:
{"named entities": [{"name": "ent_name_1", "type": "ent_type_1"}, ..., {"name": "ent_name_n", "type": "ent_type_n"}]}
Output Format:
Candidate Token: token_1
Training Data Stats: ... entity ... context ... regular ...
Contextual Meaning: ...
Relation to Examples Provided: ... positive examples ... negative examples ...
Rationale: ...
Updates: ... (add a new entity/no change)
Candidate Token: token_n
Training Data Stats: ... entity ... context ... regular ...
Contextual Meaning: ...
Relation to Examples Provided: ... positive examples ... negative examples ...
Rationale: ...
Updates: ... (add a new entity/no change)
Final predicted entities for the input text (JSON format):
Reflect False Negative - Output
Candidate Token: Italian
Training Data Stats: {"num_occurrences_as_entity": 35, "num_occurrences_as_context_tokens": 0, "num_occurrences_as_other_tokens": 0, "entity_vs_context_count": "35 vs 0", "entity_vs_non_entity_count": "35 vs 0"}
Contextual Meaning: Refers to something related to Italy, in this case, the Italian Serie A football league.
Relation to Examples Provided: Positive examples indicate that "Italian" is recognized as part of an entity (MISC) in contexts related to football or culture. There are no hard negative examples that contradict this.
Rationale: The statistical data strongly supports that "Italian" is consistently recognized as an entity, specifically as a miscellaneous term related to the context of sports.
Updates: No change
Final predicted entities for the input text (JSON format):
{"named entities": [{"name": "Italian", "type": "MISC"}]}
Reflect Boundary - Input
<input_text>
Xinhua did not say when Qinglan port in Wenchang city would be opened to foreign vessels .
</input_text>
<predicted_entity>
{"name": "Wenchang city", "type": "LOC"}
</predicted_entity>
<boundary_tokens>
[’city’]
</boundary_tokens>
<boundary_token>
city
</boundary_token>
<status>
part of the entity
</status>
<token_stat>
{"num_occurrences_as_entity": 0, "num_occurrences_as_context_tokens": 44, "num_occurrences_as_other_tokens": 20, "entity_vs_context_count": "0 vs 44", "entity_vs_non_entity_count": "0 vs 64"}
</token_stat>
<examples>
Positive Examples (part of entity):
Hard Negative Examples (context tokens):
Input: in the English city of
Output: {"name": "English", "type": "MISC"}
Negative Examples (regular tokens, neither entity nor context):
Input: said the city council would
Output: {}
</examples>
Please follow the instructions below:
1. Calibrate the boundary of the predicted entity by evaluating each boundary token listed above in relation to the predicted entity. Dont consider it if it belongs to adjacent entities. Check both the provided positive and negative examples, with particular attention to the context surrounding the boundary token and overall statistical data on inclusion or exclusion from the entity. Hard negative examples highlight tokens that are not part of the entity but are located near it.
2. Review both sets carefully. In many cases, a token may be identified as part of an entity in positive examples but not in negative ones, likely due to inconsistencies in the annotation process. if positive and (hard) negative examples seem similar, **base your decision on statistical data**, such as the frequency of the span being recognized as an entity versus its context, particularly when the data is clear-cut (e.g., one frequency is significantly higher).
3. If the token has not been seen or is rarely seen in the training data, use your best judgment to determine whether it should be considered as part of or the entire name of a **specific** entity.
4. Determine whether any modifications, such as adding or removing boundary tokens, are necessary. If changes are required, provide the updated entity by **extracting the exact text span in the sentence**, including any spaces and no outside tokens added, in JSON format. Note that abbreviations and full names are separate entities. If no tokens are added to or removed from the predicted entity, output the original entity. If all original tokens are removed, output {}.
JSON Template:
{"name": "ent_name", "type": "ent_type"}
Output Format:
Boundary Token: token_1
Training Data Stats: ... entity ... context ... regular ...
Contextual Meaning: ...
Rationale: ... positive examples ... negative examples ... data stats ...
Boundary Token: token_n
Training Data Stats: ... entity ... context ... regular ...
Contextual Meaning: ...
Rationale: ... positive examples ... negative examples ... data stats ...
Updated Predicted Entity (JSON format):
Reflect Boundary - Output
Boundary Token: city
Training Data Stats: 0 entity, 44 context, 20 regular
Contextual Meaning: The term "city" is used to describe a type of location but does not specify a unique entity on its own.
Rationale: In the provided examples, "city" appears in contexts that do not contribute to identifying a specific entity. The positive examples do not include "city" as part of the entity, while the hard negative examples show that "city" is used in a general context. The statistical data indicates that "city" has not been recognized as part of an entity in the training data.
Updated Predicted Entity (JSON format):
{"name": "Wenchang", "type": "LOC"}
Kate Deer
ICL ICL Error Reflection
Dataset Emb. Token Unseen FN Bound.
NCBI-disease 80.0 82.7 83.0 83.0 84.8
bc2gm 73.6 74.3 74.3 74.4 75.4
CoNLL03 92.9 93.5 93.5 93.5 93.6
OntoNotes 80.7 83.5 83.4 83.6 85.2
TweetNER7 62.2 63.6 63.8 63.8 64.3
Table 7: Performance breakdown of Deer (with GPT-4o) by each step in its pipeline. The subsequent reflection steps demonstrate performance gains, validating the effectiveness of the proposed reflection strategy.
Unseen Ent.
Dataset Context Seen Ent. Seen Tok. Unseen Tok.
NCBI-disease 0 90.4 59.8 70.5
1 90.4 62.1 73.3
2 90.6 64.1 76.8
3 90.8 65.4 73.8
CoNLL03 0 95.5 84.9 91.2
1 95.6 84.8 91.7
2 95.7 86.9 92.0
3 95.2 86.2 92.1
OntoNotes 0 88.1 63.7 76.2
1 89.5 67.5 80.2
2 88.6 67.8 79.6
3 89.4 68.3 78.9
Table 8: Performance breakdown of Deer on seen and unseen entities with varying context lengths (where 0 indicates the removal of context tokens). The results demonstrate that context tokens significantly enhance Deer’s performance, particularly on unseen entities.
The hyper-parameters used for Deer are presented in Table 9, and the hyper-parameters used for fine-tuning BERT-base are presented in Table 10.
Dataset $\lambda_{1}$ $\lambda_{2}$ $C$ $w_{e}$ $w_{c}$ $w_{o}$ $\theta_{FN}$ $M$ $K$
NCBI 1 1 2 1.0 1.0 0.01 0.95 1 2
bc2gm 1 0.01 2 1.0 0.5 0.01 0.9 1 2
CoNLL03 0.01 1 2 1.0 1.0 0.01 0.95 1 2
OntoNotes 1 1 2 1.0 1.0 0.01 0.95 4 2
TweetNER7 1 1 2 1.0 1.0 0.01 0.95 1 2
Table 9: Hyper-parameters for Deer. $\lambda_{1}$ and $\lambda_{2}$ control the relative contribution of two components in the retrieval. $C$ refers to context length. $w_{e}$, $w_{c}$, and $w_{o}$ weight the relative importance of three token types. $\theta_{FN}$ is set as a threshold for reflections on false negative tokens. $M$ is the number of span-level demonstrations in the reflection prompt. $K$ is the number of tokens for boundary reflection.
learning rate $1\times 10^{-5}$
# epoch 10
batch size 16
gradient accumulation steps 2
Table 10: Hyper-parameters for BERT-base fine-tuning.
Dataset Unseen FN Bound.
NCBI 92 20 130
bc2gm 23 4 163
CoNLL03 38 12 29
OntoNotes 35 21 115
TweetNER7 174 6 215
Table 11: The number of reflections triggered across five datasets using GPT-4o. The majority of reflections are concentrated on boundary tokens, while the number of reflections involving unseen tokens remains relatively small and computationally manageable.
A.2 Statistical Test on Sampled Datasets
To ensure the sampled subset provides a reliable estimate of model performance, we calculate 95% confidence intervals for Deer’s F1 scores using 1,000 bootstrap resamples. The margins of error for the three datasets are shown in Table 12 and fall within a 0.01-0.03 range, suggesting that the random sampling does not significantly affect the validity of our performance estimates.
Dataset Error Margin
bc2gm 0.027
CoNLL03 0.013
OntoNotes 0.019
Table 12: Margins of error for three sampled datasets.
Appendix B Ablation Studies
In addition to the ablation studies presented in §4.1, we further investigate the impact of other components of Deer below.
Tagging Python func. JSON
Dataset LLM (Paolini et al., 2021) (Li et al., 2023) (ours)
NCBI GPT-4o-mini 71.3 74.8 77.6
GPT-4o 80.1 81.7 82.7
Table 13: Comparing our JSON output format with prompting formats from recent works (Paolini et al., 2021; Li et al., 2023) for the ICL step in Deer, using GPT-4o-mini and GPT-4o. Our JSON format consistently outperforms the other two prompting formats, though the performance gap narrows when using GPT-4o.
B.1 Prompting Template
We compare our JSON output format (Dunn et al., 2022; Lu et al., 2022; Bai et al., 2024) with other prompting formats explored in previous studies, including tagging-based formats (Paolini et al., 2021; Wang et al., 2023a), such as “[Barack Obama | PERSON] was …”, and Python function generation (Li et al., 2023). Results in Table 13 show that our selected format outperforms two other formats, though the performance gap narrows when using GPT-4o.
B.2 Retrieval Embeddings
Following previous work (Monajatipoor et al., 2024), we experiment with seven embedding models to identify the most suitable sentence embeddings for Kate. These models include six open-source models available on HuggingFace [^7] https://huggingface.co/models via the sentence-transformer library [^8] https://sbert.net/, along with OpenAI’s proprietary model text-embedding-3-small. For open-source models, we include both general-domain models, such as all-mpnet-base-v2, and domain-specific models, such as BioClinicalBERT. Table 14 presents the performance of these seven embedding models with Kate using GPT-4o-mini. The results show that BioClinicalBERT outperforms other open-source models, highlighting the benefits of in-domain pre-training. Among all models, text-embedding-3-small achieves the best performance, surpassing others by a clear margin. Consequently, we adopt text-embedding-3-small for all datasets in our experiments.
Emb. Model Kate
bert-base-cased 70.9
bert-base-uncased 72.8
all-MiniLM-L6-v2 73.2
all-mpnet-base-v2 73.3
BioBERT 72.2
BioClinicalBERT 73.6
text-embedding-3-small 76.0
Table 14: Comparing seven embedding models to select the most suitable sentence embeddings for Kate using GPT-4o-mini. text-embedding-3-small surpasses other models by a clear margin, thus adopted for all our experiments.
B.3 Contextualized v.s. Uncontextualized Embeddings
We compare two types of token embeddings for our token-focused retriever. Contextualized token embeddings are obtained by averaging sub-token embeddings from the last layer of the embedding model, capturing sentence-level contextual information. Uncontextualized embeddings encode each token in the vocabulary independently. We evaluate both types using seven embedding models. Note that OpenAI’s text-embedding-3-small does not provide sub-token embeddings from input text, so we use it only for uncontextualized token embeddings. The generated token embeddings are then weighted by training set token-level statistics to produce sentence-level embeddings for demonstration retrieval, as introduced in §2. The results show no consistent pattern among open-source models regarding which type of token embeddings performs best; the outcome varies depending on the model. However, among all models, the uncontextualized embeddings from text-embedding-3-small achieve the best performance.
Emb. Model Contex. Uncontex.
bert-base-cased 74.9 74.9
bert-base-uncased 73.9 75.9
all-MiniLM-L6-v2 74.5 75.1
all-mpnet-base-v2 73.4 74.7
BioBERT 74.1 74.6
BioClinicalBERT 75.9 74.7
text-embedding-3-small - 76.3
Table 15: Comparing contextualized and uncontextualized token embeddings used in Deer on NCBI. No consistent pattern among open-source models is found regarding which type of token embeddings performs best. Uncontextualized embeddings from text-embedding-3-small achieve the best performance.
B.4 Two Components in Label-guided Retriever
We compare four retrieval methods: sentence embeddings used in Kate (S-Emb), our token-based retriever, and its two components, weighted token matching (T-Match) and weighted token embeddings (T-Emb). The results across three datasets are presented in Table 16. The results show that T-Emb consistently outperforms S-Emb, underscoring the importance of prioritizing entity- and context-related tokens in demonstration retrieval for NER. Combining T-Emb with T-Match improves performance, though T-Match alone lags behind S-Emb on CoNLL03. Further analysis reveals that this issue stems from short, non-language data in the dataset, such as news bylines (e.g., “OMAHA 1996-12-06”) or sports score tables, as noted by Liu and Ritter (2023). These examples contain a high proportion of unseen tokens, reducing the effectiveness of T-Match, which relies on exact token matching.
Kate Deer
Retriever S-Emb T-Match T-Emb Combined
NCBI 80.0 82.0 81.4 82.7
CoNLL03 92.9 92.2 93.3 93.5
OntoNotes 80.7 82.6 82.3 83.5
Table 16: Comparison of four demonstration retrievers: sentence embeddings (S-Emb), two components of Deer (T-Emb and T-Match), and their combination. T-Emb consistently outperforms S-Emb, highlighting the importance of prioritizing entity- and context-related tokens in retrieval. Combining T-Emb with T-Match further enhances performance.
B.5 Token Type Weights
As explained in §2, we use hyperparameters $w_{e}$, $w_{c}$, and $w_{o}$ to control the relative importance of the three token types. Their values are determined through a grid search over three predefined values: $[1.0,0.5,0.01]$. Due to budget constraints, we select $w_{e}=1.0$, $w_{c}=1.0$, and $w_{o}=0.01$ based on initial trials. This choice aligns with the intuition that entity- and context-related tokens are more important for NER. We find that these values perform well across most datasets. To evaluate the importance of each token type, we examine three sets of values, assigning $0.01$ to each weight in turn while keeping the others at $1.0$. Table 17 presents the results of the ICL step in Deer under these three settings. The results show that the default set $w_{e}=1.0$, $w_{c}=1.0$, and $w_{o}=0.01$ performs best, while the other two sets exhibit significant performance drops. This outcome verifies the critical importance of entity- and context-related tokens for NER performance.
$w_{e}$ $w_{c}$ $w_{o}$ Deer
0.01 1.0 1.0 71.6
1.0 0.01 1.0 74.7
1.0 1.0 0.01 77.6
Table 17: Comparing three sets of values, assigning $0.01$ to each weight in turn while keeping the others at $1.0$, to evaluate the importance of each token type. The default set $w_{e}=1.0$, $w_{c}=1.0$, and $w_{o}=0.01$ performs best, verifying the importance of entity- and context-related tokens for NER performance.
Figure 4: NER error ontology. This ontology is developed based on the standard strict mention-level matching metric for NER (Segura-Bedmar et al., 2013), where a predicted entity is considered correct only if both the predicted span boundaries and the entity type match with the gold ones. “Span Errors” refer to predictions that include at least span errors, while “Type Errors” denote cases where only the entity type is incorrect. “Multi-Span Errors” describe cases where multiple span issues occur, such as when a single gold entity is split into two predicted entities or when two gold entities are merged into one predicted entity. Similar analyses can be found in recent works (Ding et al., 2024; Lu et al., 2024).
Kate (GPT-4o) Deer (GPT-4o) BERT-base
Type Span Type Span Type Span
Multiple Single Multiple Single Multiple Single
Unique Paired Unique Paired Unique Paired
Dataset FP/FN FP FN FP FN FP/FN FP/FN FP FN FP FN FP/FN FP/FN FP FN FP FN FP/FN
NCBI (2014) 0 15 8 57 121 87 0 14 8 90 39 73 0 18 11 90 53 72
bc2gm (2008) 0 48 77 90 115 160 0 59 89 154 48 130 0 98 75 103 63 97
CONLL03 (2003) 53 4 8 31 44 21 54 3 6 40 23 23 66 17 16 50 28 38
OntoNotes (2013) 33 27 30 46 73 130 30 31 44 54 27 158 27 43 30 55 44 61
TweetNER7 (2022) 210 64 44 191 394 123 206 66 50 195 364 120 206 136 71 277 355 135
Table 18: Error breakdown for Kate, Deer, and BERT-base based on the hierarchical ontology of NER errors (see Figure 4). The primary source of errors across all five datasets is span-related issues for single entities. Deer, as intended by its design, performs better in reducing unique false negatives and paired errors compared to Kate. On the two datasets where Deer lags behind BERT-base, the latter still exhibits superior boundary handling, suggesting the need for a better approach to further improve span prediction.
B.6 Llama3.1-8B Inadequacy
In early experiments with Llama3.1-8B, we observed frequent failures to generate outputs in the required JSON format during ICL inference. Table 19 reports the percentage of test examples with format errors across five datasets. The failure rates range from 26.8% to 82.7%, making the model’s outputs unreliable and difficult to evaluate. Due to these inconsistencies, we exclude Llama3.1-8B from our main experiments.
Dataset Failure Rate (%)
NCBI 30.7
bc2gm 26.8
CoNLL03 81.3
OntoNotes 82.7
TweetNER7 67.5
Table 19: Percentage of example outputs from Llama3.1-8B that failed to follow the required JSON format specified in the ICL prompt.
Appendix C Error Analyses
We conduct an error breakdown based on the hierarchical ontology of NER errors in Figure 4 for Kate, Deer, and BERT-base. The full results are presented in Table 18, revealing that the primary source of errors across all five datasets is span-related issues for single entities, as highlighted in §2 when motivating the error reflection process. These errors consist of unique false positives and negatives, as well as paired errors, typically caused by boundary mismatches. Deer outperforms Kate in reducing unique false negatives and paired errors, as intended by its design.